What are the dimensions of a and b in van der waal’s equation?

Grade:12

1 Answers

piyush garg
14 years ago

vander wall equation is (p+an^2/V^2)(v-nb)=nrt

according to dimensional analysis an^2 has same dimensions as that of pv^2

hence dimensions of a=Ml^2T^-2

dimension of b is same as volume

b=(L^3)